A student cut a leaf into rectangles 50mm x 30mm. Measurements were made using a ruler (1mm graduations). The student calculated the area of the leaf. What is the percentage uncertainty for the calculated area?
1
Expert's answer
2019-10-15T14:02:10-0400

Solution:

Percentage uncertainty for the calculated area:

"\epsilon=\frac{\Delta{}A}{A}100\%"

Area of the leaf:

"A=a\cdot{}b"

Absolute uncertainty:

"\Delta{}A\approx{}dA=A'_ada+A'_bdb=b\Delta{}a+a\Delta{}b"

Then:

"\epsilon=\frac{b\Delta{}a+a\Delta{}b}{ab}100\%=(\frac{\Delta{}a}{a}+\frac{\Delta{}b}{b})100\%"

Measurements were made using a ruler with 1mm graduations. Therefore:

"\Delta{}a=\Delta{}b=0.5 mm."

Let's calculate:

"\epsilon=(\frac{0.5}{50}+\frac{0.5}{30})100\%\approx3%"


Answer: 3%